Find a if sinA= 0.82, cosA = 0.57, tanA = 1.43 and c= 60 Look at the triangle side ways or tilt head to right
B
| \
| \
| \
a | \ c
| \
| \
|————\
C b A

Answer by Alan3354(69443)

Find a if sinA= 0.82, cosA = 0.57, tanA = 1.43 and c= 60
-----------
    B
   |\   
   | \ 
   |  \
 a |   \ c = 60
   |    \
   |     \
   |______\A
       b      

sin(A) = a/c = a/60 = 0.82
a = 60*0.82 = 49.2
===========================
PS I had to stand on my head.
Tilting to the right didn't work.